From nch@roe.ac.uk Fri Apr 16 17:17:40 2010 Date: Fri, 16 Apr 2010 17:08:18 +0100 (BST) From: Nigel Hambly To: WFCAM Science Archive Team -- Eckhard Sutorius , Mike Read , Mark Holliman , Nigel Hambly , Nicholas Cross , Rob Blake , Ross Collins Cc: CCs for WSA weekly meeting minutes distribution -- Andrew Lawrence , Andy Adamson , Jim Emerson , Keith Noddle , Lorenzo Rimoldini , Mike Irwin , Peredur Williams , Bob Mann , Stephen Warren Subject: WFAU VDFS Science Archive weekly project meeting minutes, 16/04/10 Minutes of WFAU VDFS Science Archive meeting: April 16th 2010 ------------------------------------------------------------------------- ------------------------------------------------------------------------- Present: NCH, RSC, RPB, MAR, NJC, MSH, ETWS Apologies: JPE, AL, KTN, RGM DoNM: 10am, Friday April 30th 2010 in the VISTA Hut Actions discharged this week: ----------------------------- ACTION: RPB to append U/06B/H54, U/07A/H16, U/07B/H60, and U/07B/H60B to WFCAMOPENTIME for world use. Discharged; RSC reminded RPB that the Twiki ops note needs updating to reflect this. Actions partly discharged but continuing: ----------------------------------------- None this week. Actions carried forward from 09/04/10 meeting: ---------------------------------------------- ACTION: NCH to fill out a Q2 2010 plan and continue with the progress monitoring for the time being. continues Specific points and new actions: -------------------------------- Project management: NCH noted that VSA requirements are becoming slowly clearer following a few discussions with a few PIs at NAM. WFCAM & VISTA updates: Nothing of note this week. Comments and issues arising from CASU minutes: No new minutes this week, but it was noted that WFCAM 10A data is piling up at the CASU end with no ok-to-copy flags releasing it to WFAU. This is a little puzzling since, for example, a Japanese non-survey user contacted the WSA in the week for their 10A data, and when we couldn't oblige CASU kindly volunteered to send them their data. But hang on... if it's OK for the PI to copy, shouldn't it be OK for WFAU to copy?! If we didn't know better we might worry that CASU don't love us anymore ... Networking: Nothing of note this week. WSA/VSA Operations: ETWS noted "- Ingest of VISTA data is progressing with some minor hickups: - Some files showed a value of infinity in halfFluxErr, I'll send a full list to CASU when processing has finished. - A new Programme (284.A-5026(A)) showed up. According to the ESO query page it's doing early science with the VISTA NB118 filter. Otherwise all the metadata up to 31/01/10 is ingested (CU3) and ingest of the catalogue data (CU4) for the SV period (15/10/09 - 02/11/09) has nearly finished. The DRYRUN data will be ingested over the weekend. Creation of JPGs is also going along smoothly and might finish mid of next week. - We are still waiting for the release of the last bit (02/10) of v0.8 VISTA data to finish the DRYRUN period. - I also started copying the reprocessed UDS J-band data." ETWS and RPB noted that the version 0.8 Vista data is nearly ready for release (see above) so flat-file and release DBs incorporating the fully calibrated detection data in DB tables will be released to the PIs. Enhanced DB-driven products will be included for the wide-shallow surveys (VHS/VIKING) where possible; while first-pass auto-QC and generation of the DB-driven deep products will be done asap in a second release. ACTION: MAR to make a first-pass auto QC script for broad-brush crapulence filtering of VSA data ACTION: NJC to continue working closely with JRL at CASU on implementing the archive-end pixel processing codes. Hardware and Systems: Further to last week's minutes, it seems that the sub-optimal RAID filesystem disk05 is knackered. RSC noted that this is the disk used as temporary workspace by CUs 7+16, so ACTION: RSC to reconfigure CUs that use disk05 as temporary storage space to use the nifty disk34 instead. MSH asked to whom orders for new equipment should be given for approval. Looks like RGM will have a queue outside his door when he returns from Rio on Monday (volcanic activity permitting). Software: NJC is continuing to modify CUs 13 and 14 inline with changes to the pixel processing and source extraction tools from CASU; RSC will be working on the implementation of proper motion determination into the CU7 source merging software. MAR noted that there's a new name-resolver in the UI that will now appear in webforms that prompt for coords; and also that the VSA download limit will be increased from 0.2 to 1GB following a request from RGMcM at IoA. (MSH noted that VO services have no such download limit). Survey Data Release: It was noted that GPS DR4 will go world-public next week (on April 22nd). ACTION: RPB to switch on world read permissions on all GPS DB objects in UKIDSSDR4PLUS. Non-survey Data Release: RPB noted that non-survey catalogue DBs for 09A/B projects continue to be released to their respective PIs. Astrogrid deployment & Data Analysis services: MSH noted that the move to new web server seshat caused a few hiccups, in particular the SIA services were briefly broken, but these are now fixed. Miscellaneous: Nothing else this week. ============================================================= Nigel Hambly Tel: +44-131-668-8234 Institute for Astronomy Fax: +44-131-668-8416 School of Physics and Astronomy University of Edinburgh Email: nch@roe.ac.uk Royal Observatory Blackford Hill Edinburgh EH9 3HJ The University of Edinburgh is a charitable body, registered in Scotland, with registration number SC005336. =============================================================